Home‎ > ‎Crew 133‎ > ‎

112-CommanderReport

Commander Report
01/12/2014
Paula Crock
 
We are at the mid-way point of our rotation.  I’m a little too tired to do the math to figure out
whether we’re into the second half yet but I think we are.  Days go by so fast it feels like a month
ago that we left Earth.  I found a ten dollar bill in my pocket today and it looked strange to me.
When I commented about that to Gartrelle he said this is the longest in his adult life he’s gone
without spending money.  Life really is different out here.  No traffic, no stress of evening plans,
no grocery shopping, no phone calls.  We’re acclimated to it now, so that ten dollar bills seem
strange.  I think the first few ring tones we hear will be strange too, as will turning on a water
faucet and not hearing the thump-thump-thump of the water pump.   But we still have a week before we
go back to that world.
 
Today we had another EVA -- data collection for Jessup’s strata study and my helmet temperature and
radio comms studies.  The weather was rough today but the geology was unworldly and worth fighting
the wind to see.  Towards the end of the EVA Diaz-Rubin and I did some extra geology exploration
while Jessup and Komorowski hiked up a hill above us to survey the surrounding area better.  Upon
return to the Hab we began the surprise birthday preparations for Howell while Gartrelle distracted
her with hydroponics prep in the GreenHab.  Komorowski’s cake was a treat!  I hope Howell’s birthday
at MDRS was a memorable one.

-----------------------

CommanderCheckIn Report
01/12/2014
Paula Crock

Crew Physical Health:
All are well.
  
Eva Depart Time:
10:02
Eva Return Time:
12:15
Narrative Of Field Mission Results:
Collected data on two strata points for Jessup study and 5 radio check-in
tests for Crock study
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Eva Depart Time:
 
Eva Return Time:
 
Narrative Of Field Mission Results:
 
Hab Maint:
EVA helmet repair
Bathroom vent fan investigation
Rep Sched:
Commander Check-In
Engineering Report
Science Report (GreenHab)
Science Report (Geology)
Science Report (Helmet Temp and Radio Signal)
EVA Report
Daily Photos
Journalist Report
Commander Report
Tomorrow Plan:
Another all in-Hab day for projects; possibly starting hydroponics if we
have sufficient water
Inventory:
N/A
Support Req:
Still awaiting word on whether there are replacement bulbs/ballasts on
station for the lab lights
Miscelleneous:
N/A